### v1

# MDPE: A Multimodal Deception Dataset with Personality and Emotional Characteristics

Abstract
Deception detection has garnered increasing attention in recent years due to the significant growth of digital media and heightened ethical and security concerns. It has been extensively studied using multimodal methods, including video, audio, and text. In addition, individual differences in deception production and detection are believed to play a crucial role.Although some studies have utilized individual information such as personality traits to enhance the performance of deception detection, current systems remain limited, partly due to a lack of sufficient datasets for evaluating performance. To address this issue, we introduce a multimodal deception dataset MDPE  [^1] The dataset is available at https://github.com/cai-cong/MDPE, licensed under CC BY-NC-SA-4.0 license.. Besides deception features, this dataset also includes individual differences information in personality and emotional expression characteristics. It can explore the impact of individual differences on deception behavior. It comprises over 104 hours of deception and emotional videos from 193 subjects. Furthermore, we conducted numerous experiments to provide valuable insights for future deception detection research. MDPE not only supports deception detection, but also provides conditions for tasks such as personality recognition and emotion recognition, and can even study the relationships between them. We believe that MDPE will become a valuable resource for promoting research in the field of affective computing.
1 Introduction
Generally, deception refers to the act of misleading, tricking, or deceiving others [1]. It involves hiding the truth or presenting false information to create an impression that is not accurate. Deception can take many forms, including both verbal and nonverbal information [2]. And it also occurs in various contexts, such as interpersonal relationships, business, politics, and entertainment. Deception is often considered unethical and can have serious consequences for trust and relationships.
As deception has expanded to other fields such as social media, interviews, online transactions, and deception in daily life,the need arises for a reliable and efficient system to aid the task of detecting deceptive behavior. Many machine learning approaches have been proposed in order to improve the reliability of deception detection systems [3]. In particular, physiological, psychological, visual, linguistic, acoustic, and thermal modalities have been analyzed in order to detect discriminative features and clues to identify deceptive behavior [4, 5, 6, 7]. Video-based deception detection is a current priority in deception research, because behavioral cues can be extracted from videos in a cheaper, faster, and non-invasive manner [8], which is preferable to invasive approaches that extract clues through devices attached to human bodies (e.g., polygraphs). Visual clues of deception include facial emotions, expression intensity, hands and body movements, and microexpressions. These features were shown to be capable of discriminating between deceptive and truthful behavior [9, 10]. Acoustic features took into account the pitch and speaking rate among other measurements to specify whether certain features are associated with an act of deceit [5]. Linguistic features were usually extracted from the language, words usage, and consistency of the statements made by a person [11]. Recently, multimodal analysis has gained a lot of attention due to their superior performance compared to the use of unimodal modalities. In the deception detection field, several multimodal approaches [12, 13, 14, 15] have been suggested to improve deception detection by integrating features from different modalities. This integration created a more reliable system that is not susceptible to factors affecting sole modalities and polygraph tests.
In addition, it is firmly believed that there are individual differences in deception production and detection [16, 17, 18]. Specifically, it includes cognitive level, personality traits, psychological characteristics, and emotional expression. Everyone has different personalities and psychological characteristics, and the expression of emotions is also various. It has been demonstrated through several studies that personality factors and emotional cues play a significant role in subjects’ ability to deceive and detect deception [16, 19]. Emotion is a fundamental aspect of human communication that interacts with cognition, guiding social behavior in both human-to-human interactions and human-computer interactions [20, 21]. Emotional characteristics are important, because deceptive behavior can trigger emotional states, leading to behavioral changes that serve as deceptive clues [9, 22]. However, it is difficult to directly improve the accuracy of deception detection using emotional features [23]. One of the reasons is that emotional expression is also a part of deception. It is usually difficult to detect whether a deceiver’s emotional expression is genuine or disguised.
To address this issue, we propose a multimodal deception dataset MDPE. It not only collects subjects’ deception information, but also personality information and emotional expression information. Each subject was required to conduct another emotional experiment in addition to engaging in deception, in order to obtain their true emotional expression. Although our research was conducted in the laboratory to provide clear and comparable conversations, we provided subjects with effective monetary incentives to detect and generate effective deceptive behavior [16]. To our knowledge, this is the largest multimodal deception dataset in the released dataset and the only deception detection dataset with personality and emotional characteristics.
To sum up, our contributions are threefold:
  • We propose a novel multimodal deception dataset MDPE with personality and emotional characteristics, composed of facial video, and audio recordings and transcript. And an easily replicable experimental protocol has also been provided to researchers.
  • We provide a benchmark for deception detection from multimodal signals, and discussed the impact of personality traits and emotional cues on deception detection.
  • We offer new possibilities to facilitate further affective computing research, encourage the development of new methods that utilize individual differences for deception detection, as well as for tasks such as personality recognition and emotion recognition.
2 Related Work
Deception Dataset   Pérez-Rosas et al. [12] introduced a new multi-modal deception dataset Real-life Trial having real-life videos of courtroom trials. They demonstrated the use of features from different modalities and the importance of each modality in detecting deception. They also evaluated the performance of humans in deception detection and compared it with their machine learning models. The Box-of-Lies dataset [24] was released with video and audio from a game show, and presents preliminary findings using linguistic, dialog, and visual features. Multiple modalities have been introduced in the hope of enabling more robust detection. Pérez-Rosas et al. [25] introduced a dataset for deception including video and thermal imaging, as well as physiological and audio recordings. Gupta et al. [26] proposed Bag-of-Lies, a multimodal dataset with gaze data for detecting deception in casual settings. Speth Jeremy et al.[27] proposed a multimodal deception database DDPM contains almost 13 hours of recordings of 70 subjects, as well as physiological signals such as thermal video frames and pulse oximeter data. Most studies on deception detection are designed and evaluated on private datasets, typically with relatively small sample sizes, and MDPE dataset addresses these drawbacks. Table 1 compares the sample size and length for existing datasets and MDPE.
Table 1: Comparison of the subject count and length for several databases for deception detection
dataset Subjeet Count Length(Minutes)
Multimodal 30 -
Real Trials 56 56
Box-of-Lics 26 144
Bag-of-Lies 35 <241
DDPM 70 776
MDPE 193 6209
Multimodal Deception Detection   Decades of research in psychology, and deception detection have documented verbal and nonverbal behavioral cues indicative of deceptive communication. Visual cues such as the frequency and duration of eye blinks [28, 29, 30], dilation of pupils [31, 32], and facial muscle movements [33, 34] have been found to distinguish between deceptive and truthful behavior. Vocal cues can be indicative of deception, with deceptive speakers tending to speak with higher and more varied pitch [1, 35], shorter utterances, and less fluency [36, 37] than truthful speakers. Deception also correlates with verbal attributes of speech, with deceivers tending to communicate with less cognitive complexity, fewer self-references, and more words indicative of negative emotions [38, 6]. Mohamed Abouelenien et al. [39] explored a multimodal deception detection approach and integrates multiple physiological, linguistic, and thermal features. They used a decision tree model, to gain insights into the features that are most effective in detecting deceit. Leena Mathur et al. [15] analyzed the discriminative power of features from visual, vocal, and verbal modalities affect for deception detection. They experimented with unimodal Support Vector Machines (SVM) and SVM-based multimodal fusion methods to identify effective features for detecting deception.
Individual Difference Deception   Some studies confirm that some of the five NEO-FFI (Neuroticism-Extraversion-Openness Five-Factor Inventory) dimensions are related to deception [40, 41]. Sarah Ita Levitan et al. [16] reported the role of personality factors derived from the NEO-FFI and of gender, ethnicity and confidence ratings on subjects’ ability to deceive and to detect deception. Justyna Sarzyńska et al. [42] reports correlations between the ability to lie and extraversion, as well as conscientiousness. Personality characteristics are a promising set of information for deception detection, and similarly, emotional characteristics are also important. Joseph P. Gaspar et al. [43] integrate prior theory and research on emotions, emotional intelligence, and deception and introduce a theoretical model. This model explores the interplay between emotional intelligence (the ability to perceive emotions, use emotions, understand emotions, and regulate emotions; and deception. Mircea Zloteanu et al hold strong beliefs about the role of emotional cues in detecting deception, and explored how decoders’ emotion recognition ability and senders’ emotions influence veracity judgements [44]. Joseph P. Gaspar et al. [19] believe that emotions are both an antecedent and a consequence of deception, and they introduce the emotion deception model to represent these relationships. This model broadens their understanding of deception in negotiations and accounts for the important role of emotions in the deception decision process. To our knowledge, MDPE is the only deception detection dataset with personality and emotional characteristics.
3 Dataset
3.1 Materials
We collect our deception dataset using: a sports camera Gopro Hero9 with a resolution of 1920x1080 and a frame rate of 60 fps. The voices of the subjects are also recorded by the built-in microphone of the camera. A Thinkpad laptop was provided to subjects for watching emotion induction videos during the emotion experiment. The experimental place is in a professional recording stdio, and during the data collection process, only the subject and interviewer stay in the room. Some materials were prepared by the Data Collection Coordinator (DCC). The experimental setup is shown in Figure 1.
Figure 1: The example of the place setup for data acquisition.
3.2 Participants
There were 193 subjects in this study, of which 130 were female and 63 were male. They are all native Chinese speakers from different backgrounds. Their age ranged from 18 to 69 years old, and they had various professions including students, workers, teachers, retirees, etc.
Firstly, we segmented the raw video, resulting in 1808 minutes of deceptive video and 4401 minutes of emotional video, totaling 6209 minutes. Each of the 193 subjects provided 24 responses, 9 of which were deceptive. The length of deception videos ranged from 4 minutes to 27 minutes. Each subject had 16 emotional videos, with lengths ranging from 19 minutes to 38 minutes (including the time spent watching emotional induction videos).
Before the experiment began, the subjects were informed of all experimental procedures. The subjects explicitly consented to record their conversation and publish the video data in a scientific conference or journal. And we do not publish any privacy sensitive data, and the anonymity of participants will be guaranteed. All data were collected under a protocol approved by the authors’institution’s Human Subjects Institutional Review Board.
3.3 Procedure
Personality Characteristics Collection: Subjects were required to fill out a Big Five personality questionnaire [45], which consists of 60 questions. Each question was marked with a score indicating whether the descriptions match their own, with 1 indicating strong disagreement and 5 indicating strong agreement. Details can be found in the appendix A.
Emotional Experiment: Subjects were asked to watch 16 emotional induction videos, including two induction videos for each emotion of sadness, happiness, relaxation, surprise, fear, disgust, anger, and neutral. There are a total of 39 induction videos, of which 17 are from the Chinese Emotional Video System (CEVS) [46]. Each video segment has been labeled and evaluated to ensure that it can induce corresponding emotions. Another 22 are from our online collection. Because the CEVS only include six emotions: sadness, happiness, fear, disgust, anger, and neutral, and some videos of the CEVS are outdated and cannot successfully induce corresponding emotions in our pre-experiments. Each video we collected online was annotated by 12 data annotators based on CEVS selection criteria and evaluation methods, and the results showed that each video triggered strong emotions.
Before the emotional experiment began, the DCC randomly selected 16 induction videos (ensure two videos for each emotion) for the subjects to watch. After watching each video, subjects were required to describe their feelings and then fill out an “Emotional Scale", which quantified 8 emotions. Subjects rated their 8 emotions, ranging from 1 to 5, with 1 indicating no such emotion and 5 indicating the strongest emotion. Details of the emotion scale can be found in the appendix B.
Deception Experiment: The deception data collection process follows DDPM [27]. The interviewer conducted an interview with the subject, asking a total of 24 questions. These 24 questions were jointly formulated by 5 psychology researchers with over 5 years of experience. Before the interview, details of the emotion scale can be found in the appendix C. The DDC randomly selected 9 questions that must lie and hand them over to the subject (the interviewer does not know which 9 questions). The first 3 questions will not be selected, which means that the first three “warm up” questions were always to be answered honestly. They allowed the subject to get settled, and gave the interviewer an idea of the subject’s demeanor when answering a question honestly.
The subject have a maximum of 15 minutes to prepare, and during the preparation process, they must remember these 9 questions and think about how to deceive in the upcoming interview process. During the interview process, when asked these 9 questions, the subject must lie, and when asked the remaining 15 questions, they must tell the truth. Subjects were motivated to deceive successfully through two levels of bonus compensation: if they were able to deceive the interviewer in five or six of the nine deceptive responses, they were given a 150 percent of a base incentive payment; the base payment was doubled if they were successfully deceptive in seven or more questions. In order to collect more indistinguishable deception answers, we encourage subjects to incorporate some truth into lies when answering these deceptive questions.
During the interview process, the interviewer asked 24 questions in random order, and provide their judgment of truthful or deceptive answers to each question. And the interviewer filled out the "Interviewer Judgment Scale", which record the trust level the interviewer thinks of each answer. The trust level was divided into 1-5 points, where 1 represents definitely true and 5 represents definitely false. After the interview, the subject also filled out the "Subject Lie Confidence Scale" and be asked to rate the answer they just lied to. The same score is 1-5, where 1 represents that I have definitely deceived successfully and 5 represents that I have definitely not deceived successfully.
Each subject was required to complete the above three experiments, so that we can obtain their personality , emotional and deceptive characteristics.
4 Benchmark
4.1 Data Preprocessing
For the visual modality, we first unify the raw video to 30 fps of the frame rate, and crop and align faces via DLib Toolkit [47]. Figure 2 shows some examples of real and deceptive faces. Then, we use visual encoders to extract frame-level , followed by average pooling to compress them to the video level. For the audio modality, we use FFmpeg to separate the audio from the video and unify the audio format to 16kHz and mono. For the textual modality, we first extract transcript using Paraformer [48], an open-source ASR toolkit. After data preprocessing, for each sample $x_{i}$, we extract acoustic features $f_{i}^{a}\in\mathbb{R}^{d_{a}}$, textual features $f_{i}^{l}\in\mathbb{R}^{d_{l}}$, and visual features $f_{i}^{v}\in\mathbb{R}^{d_{v}}$, where $\left\{d_{m}\right\}_{m\in\left\{a,l,v\right\}}$ is the feature dimension for each modality.
4.2 Feature extraction
Different features lead to distinct results. To guide feature selection, we evaluate the performance of different features under the same experimental setup.
Visual Modality: Compared with handcrafted features, deep features extracted from supervised models are useful for facial expression recognition [49]. CLIP [50] is a multimodal model based on contrastive learning, where training utilizes text and images to construct positive and negative sample pairs. Pre-training is conducted on a dataset comprising 400 million pairs, resulting in strong generalization capabilities. And Vision Transformer (ViT) [51] is a transformer encoder model , pre-trained in a supervised manner on a large dataset of images. Images are presented to the model in the form of sequences of fixed-size patches (resolution of 16x16) and undergo linear embedding.
Acoustic Modality: We extracted the handcrafted feature extended Geneva Minimalistic Acoustic Parameter Set (eGeMAPS) [52], which contains 88 acoustic parameters designed specially for speech emotional recognition tasks, covering spectral, cepstral, and prosodic features . And Wav2vec [53] demonstrates the power of learning robust representations solely from speech audio, fine-tuning on transcribed speech, surpassing the best semi-supervised methods. It masks speech inputs in the latent space and addresses a contrastive task defined on quantized latent representations. It has been widely applied to downstream speech tasks. HUBERT [54] utilizes offline clustering steps to provide aligned target labels for prediction losses. A key component of this approach is applying prediction losses only in masked regions, forcing the model to learn representations that combine acoustic and language models on contiguous inputs. To better distinguish between speakers, a sentence mixture training strategy WavLM [55] is proposed, allowing for the unsupervised creation and merging of additional overlapping sentences during the training process.
Text Modality: Among all language models, BERT [56] and its variants are widely utilized, and it uses the masked language model and next sentence prediction objectives to learn word embeddings. Here, we extract the sbert-chinese-general-v2 features, which is based on the bert-base-chinese version of the BERT model and is trained on the million-level semantic similarity data set SimCLUE. ChatGLM [57] is an open-source conversational language model supporting bilingual question answering in both Chinese and English, based on the General Language Model (GLM) architecture. It demonstrats exceptional contextual understanding and more efficient inference capabilities. Baichuan [58] is an open-source large-scale model with 13 billion parameters. It features a larger size, more extensive training data, and more efficient inference capabilities.
Figure 2: Some examples of truth and deception faces.
4.3 Model Structure
For unimodal features, we utilize the fully-connected layers to extract hidden representations and predict deception:
$h_{i}^{m}=\text{ReLU}\left(f_{i}^{m}W_{m}^{h}+b_{m}^{h}\right),{m\in\left\{a,l,v\right\}}$ (1)
$\hat{y}_{i}=\text{softmax}\left(h_{i}^{m}W_{m}^{d}+b_{m}^{d}\right),m\in\left\{a,l,v\right\}$ (2)
where $h_{i}^{m}\in\mathbb{R}^{h}$ is the hidden feature for each modality, $d_{i}\in\mathbb{R}^{2}$ is the estimated deception probabilities. For multimodal features, different modalities contribute differently to deception detection. Therefore, we compute importance scores $\alpha_{i}\in\mathbb{R}^{3\times 1}$ for each modality and exploit weighted fusion to obtain multimodal features:
$h_{i}=\text{Concat}\left(h_{i}^{a},h_{i}^{l},h_{i}^{v}\right)$ (3)
$\alpha_{i}=\text{softmax}\left(h_{i}^{T}W_{\alpha}+b_{\alpha}\right)$ (4)
Similarly, for personality and emotional expression features, we use concatenation for feature fusion. We directly use personality scale scores as personality features. For emotional features, we train an emotion recognition model first, input all emotional expression samples into the emotion recognition model, and take the last fully connected layer features for average pooling as the emotion expression feature.
Table 2: Unimodal results on MDPE. "P" denotes the addition of personality features,"E" denotes the addition of emotional features
Feature Accuracy AUC with P with E with P and E Accuracy AUC Accuracy AUC Accuracy AUC VIT 60.30% 0.577 61.27% 0.582 61.43% 0.615 61.55% 0.620 CLIP-base 58.54% 0.602 59.17% 0.611 58.32% 0.606 59.11% 0.605 CLIP-large 57.30% 0.574 58.34% 0.582 56.97% 0.555 57.67% 0.579 eGeMAPS 55.86% 0.563 57.22% 0.588 56.22% 0.577 56.89% 0.585 HUBERT-base 58.13% 0.615 62.38% 0.651 59.35% 0.617 62.12% 0.641 HUBERT-large 60.80% 0.636 62.07% 0.646 60.34% 0.621 61.87% 0.644 Wav2vec2-base 58.75% 0.581 59.74% 0.603 59.99% 0.594 59.84% 0.599 Wav2vec2-large 60.10% 0.582 61.88% 0.617 59.32% 0.592 62.10% 0.634 WavLM-base 61.66% 0.609 60.82% 0.606 60.16% 0.595 60.92% 0.593 WavLM-large 57.82% 0.599 60.31% 0.583 58.02% 0.607 60.52% 0.611 Sentence-BERT 61.76% 0.639 62.34% 0.651 63.21% 0.641 63.34% 0.657 ChatGLM2-6B 60.73% 0.648 61.45% 0.659 61.45% 0.667 61.56% 0.676 Baichuan-13B 61.87% 0.649 62.90% 0.667 63.32% 0.675 63.74% 0.683
4.4 Implementation Details
We select the dimension of latent representations from $\left\{64,128,256\right\}$. During training, we use the Adam [59] optimizer and choose the learning rate from $\left\{10^{-3},10^{-4}\right\}$. We set the maximum number of epochs to 300 and the weight decay to $10^{-5}$. Dropout [60] is also employed, and we select the rate from $\left\{0.2,0.3,0.4,0.5\right\}$ to alleviate the over-fitting problem. We randomly select 5 answers (3 truths and 2 deceptions) from 24 answers in all samples as the validation set, and the remaining 19 answers as the training set. To mitigate randomness, we run each experiment five times and report the average result. And we choose the cross-entropy as loss function and the accuracy and AUC as the evaluation metric.
4.5 Experiment Results
Unimodal Results In this section, we establish the unimodal benchmark for MDPE and report results in Table 2. We hope this benchmark can provide guidance for feature selection and point the way to developing powerful feature extractors. For the visual modalities, VIT achieves better results than CLIP, possibly because VIT is trained on supervised datasets and can reveal more deceptive clues than CLIP features that use text as a supervisory signal. For the acoustic modality, the deep features outperform handcrafted features. This may be because egemas are features related to emotions, using a emotion-related handcrafted acoustic feature may limit performance. In contrast, deep features can capture more universal acoustic representations for deception detection. For the textual modality, we focus on textual encoders that support Chinese (large language models are generally trained on multilingual corpora containing Chinese). And the textual modality can achieve best performance than the visual and acoustic modalities, which indicates that our dataset textual reveals more deceptive clues than other modalities.
Of course, we have also incorporated personality and emotional features into deception detection. Models trained with personality features almost often exhibit superior performance. These findings demonstrate the importance of personality in deception detection tasks. And the addition of emotional features has also improved the performance of deception detection models, but the improvement is not comparable to personality features, even some results have decreased. This may be because personality traits are directly usable features, and emotional features are extracted by emotion recognition models. The quality of features is also influenced by the emotion recognition models. In the future, better methods for using emotional expression features can be explored. By incorporating both personality and emotional features, the deception detection model achieved the highest performance in unimodal results, These results show that personality and emotional expression characteristics are indeed important for deception detection tasks, and using them can truly achieve deception detection modeling based on individual differences. It is worth mentioning that among all the unimodal results, Baichuan always achieved the best results. It is the model with the largest number of parameters among the features we used, and it further demonstrates the potential of large language models in deception detection.
Multimodal Results In Table 3, we select several well-performing unimodal features and report their fusion results. Experimental results demonstrate that multimodal fusion consistently improves performance. The reason lies in the fact that deception cues can be conveyed through multiple modalities. The integration of multimodal information allows the model to better comprehend the video content and accurately detection deception. Firstly, almost all features have achieved performance improvements in multimodal feature fusion, but the fusion of visual and acoustic modalities has hardly improved or even declined. It indicates that the addition of text features make the performance of the model tend to stabilize. This is consistent with human judgment. When people judge whether others are lying, they tend to express the truth or falsehood of the content, because it difficult to judge from visual or acoustic features. In the future, further exploration and research should be conducted on deceptive clues in visual and acoustic features.
Similar to unimodal results, models trained with personality features typically exhibit excellent performance. The addition of emotional features makes the performance of the model unstable. Some features have been improved, while others have decreased. Finally, the best result always comes from the three modalities fusion, and the three modalities fusion with personality and emotional characteristics has achieved the best performance.
Table 3: Multimodal results on MDPE. We select several well-performing unimodal features and report their fusion results. Here, “V”, “A” and “T” represent the visual, acoustic and textual modalities, respectively.
V A T Accuracy AUC with P with E with P And E Accuracy AUC Accuracy AUC Accuracy AUC VIT HBB - 61.76% 0.628 61.53% 0.617 61.14% 0.616 62.68% 0.628 VIT WMB - 60.88% 0.611 60.22% 0.599 60.02% 0.599 60.72% 0.587 CLB HBB - 61.02% 0.618 60.14% 0.606 60.53% 0.614 61.45% 0.625 VIT - Bai 63.31% 0.664 63.47% 0.678 63.31% 0.676 63.48% 0.672 CLB - Bai 63.31% 0.666 63.52% 0.677 63.38% 0.667 63.10% 0.664 CLL - Bai 64.15% 0.665 63.94% 0.679 63.98% 0.672 64.04% 0.678 - HBL Bai 62.69% 0.658 63.42% 0.665 63.00% 0.665 63.42% 0.663 - W2B Bai 63.83% 0.663 63.48% 0.663 63.57% 0.668 63.79% 0.664 - WMB Bai 64.25% 0.661 64.15% 0.679 63.90% 0.677 64.07% 0.679 VIT HBB Bai 64.45% 0.675 64.33% 0.679 63.93% 0.674 64.00% 0.675 VIT WMB Bai 63.42% 0.666 64.87% 0.681 63.62% 0.664 63.59% 0.672 CLB HBB Bai 62.94% 0.657 63.93% 0.678 63.97% 0.678 64.66% 0.687
5 Discussion
5.1 Future work
Firstly, dimensional emotions are also important. In fact, we are labeling MDPE with dimensional emotions (including deception and emotion), which can not only study the impact of dimensional emotional features on lie detection at the individual level, but also explore more emotional clues in the deception process. Secondly, we extracts deep features from some pre-trained models and uses simple models for deception detection. In the future,larger models should be designed to be used for deception detection tasks. Thirdly, this article simply concatenates personality and emotional expression features to assist in deception detection tasks. In the future, more complex feature fusion or model fusion algorithms can be used for deception detection. Finally, in this paper, we only focuses on deception detection tasks, but MDPE includes individual personality, deception, and emotional expression information, which can not only support deception detection tasks, but also tasks such as personality recognition and emotion recognition. Even these individual level information can be used to assist other tasks, as shown in Figure 3. In fact, there have been many studies on the relationship between personality and emotion [61, 62, 63], but the lack of relevant datasets has led to slow research progress in this field, and we believe that MDPE can provide valuable resources for these research directions.
Figure 3: Possible future work on the MDPE
5.2 Limitations
Firstly, although the subjects were required that they must lie about the deception questions, and verified the deceptive questions and content with the Interviewer after the deception experiment, we do not know whether the subjects have actually deceived on the deception questions. Secondly, although our induction videos have been annotated by professionals to demonstrate their reliability and validity, there are still some subject whose feelings are inconsistent with the emotions we expected to induce. This is because different people have different understandings of the video content, triggering different emotions. Thirdly, relying on self-assessment scales for data annotation is a subjective process for subjects, which may lead to bias in subsequent analysis. Different subjects may have significant differences in their perception of emotions. In addition, MDPE only collects native Chinese speakers, there may be cultural differences in deception detection. Finally, gender imbalance among subjects in MPDE is a common issue in human data collection [64, 65].
6 Conclusion
We present MDPE, a dataset for deception detection featuring three categories of data modalities, as well as personality and emotional characteristics. Firstly, diverse modalities carry complementary information that can be jointly exploited. By providing access to multiple synchronized modalities, MDPE enables cross-modal analyses that have the potential to improve the understanding of the relationships between video, audio, and text. Secondly, it can help improve the understanding of deceive behavior, aiming to develop reliable deception detection algorithms and enhance the security issues related to deception in our society. Thirdly, MDPE provides the personality traits and emotional expression characteristics of each subject, which can help analyze the impact of personality and emotional expression on deceive behavior. In addition, MDPE not only supports deception detection models, but also provides conditions for personality recognition and emotion recognition tasks, and can even study the relationship between deception, personality, and emotion, such as using personality features to improve the performance of emotion recognition tasks. Finally, to promote reproducibility, MDPE also provided a set of benchmark experiments. Although the proposed model focuses on deception detection, the use of personality and emotional features also demonstrates the predictive potential of our dataset. They represent a good starting point for future work that researchers and developers can use as a benchmark. By openly sharing MDPE, we hope to ignite new advances in this critical area of affective computing.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes] See 5.2
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] See 5.2
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [Yes]
    2. (b)
      Did you include complete proofs of all theoretical results? [Yes]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes]
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See 4.4
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [No]
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [No]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes]
    2. (b)
      Did you mention the license of the assets? [Yes]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] See 3.2
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [Yes] See 3.2
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [Yes]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes]
Appendix A Big Five Personality Inventory Second Edition (BFI-2)
Below are some descriptions of personal characteristics, some may or may not apply to you. Please fill in the corresponding number on the horizontal line before each sentence below to indicate whether you agree or disagree with this description.
  1. 1.
    Outgoing personality, enjoys socializing
  2. 2.
    Soft hearted and compassionate
  3. 3.
    Lack of organization
  4. 4.
    Calm and adept at handling pressure
  5. 5.
    Not very interested in art
  6. 6.
    Strong and confident personality, daring to express one’s own opinions
  7. 7.
    Humble and respectful towards others
  8. 8.
    Relatively lazy
  9. 9.
    Being able to maintain a positive attitude even after experiencing setbacks
  10. 10.
    Interested in many different things
  11. 11.
    I rarely feel excited or particularly want to do anything
  12. 12.
    Often picking on others’ faults
  13. 13.
    Reliable and reliable
  14. 14.
    Irregular mood and frequent emotional fluctuations
  15. 15.
    Skilled in creativity and able to find smart ways to do things
  16. 16.
    Relatively quiet
  17. 17.
    Lack of empathy towards others
  18. 18.
    Work in a planned and organized manner
  19. 19.
    Easy to get nervous
  20. 20.
    Enthusiastic with art, music, or literature
  21. 21.
    Often in a dominant position, like a leader
  22. 22.
    Often having disagreements with others
  23. 23.
    It’s difficult to start taking action to complete a task
  24. 24.
    Feeling secure and satisfied with oneself
  25. 25.
    Disliking discussions with strong knowledge or philosophy
  26. 26.
    Not as energetic as others
  27. 27.
    Be magnanimous and magnanimous
  28. 28.
    Sometimes I lack a sense of responsibility
  29. 29.
    Emotionally stable and less likely to get angry
  30. 30.
    Almost no creativity
  31. 31.
    Sometimes shy and introverted
  32. 32.
    Helpful and selfless towards others
  33. 33.
    Habit keeps things tidy and orderly
  34. 34.
    Often worried and worried about many things
  35. 35.
    Valuing Art and Aesthetics
  36. 36.
    Feeling difficult to influence others
  37. 37.
    Sometimes being rude to people
  38. 38.
    Efficiency, starting and ending with work
  39. 39.
    Often feeling sad
  40. 40.
    Deep thinking
  41. 41.
    Full of energy
  42. 42.
    Do not trust others and doubt their intentions
  43. 43.
    Reliable, always trustworthy to others
  44. 44.
    Able to control one’s emotions
  45. 45.
    Lack of imagination
  46. 46.
    Loud and talkative
  47. 47.
    Sometimes cold and indifferent to others
  48. 48.
    It’s messy and doesn’t like to tidy up
  49. 49.
    Rarely feel anxious or afraid
  50. 50.
    Feeling bored with poetry and drama
  51. 51.
    I prefer to have others take the lead and take responsibility
  52. 52.
    Humility and courtesy towards others
  53. 53.
    Have perseverance and be able to persist in completing tasks
  54. 54.
    Often feeling depressed and unhappy
  55. 55.
    Not very interested in abstract concepts and ideas
  56. 56.
    Full of enthusiasm
  57. 57.
    Think about people in the best possible way
  58. 58.
    Sometimes they may engage in irresponsible behavior
  59. 59.
    Emotions are variable and prone to anger
  60. 60.
    Creative and able to come up with new ideas
Appendix B Emotional Sacle
After watching the video, you need to rate the following emotions: sadness, relaxation, happiness, surprise, fear, anger, disgust, and neutral. Mark to what extent you feel it appropriately expresses your feelings, with intensity ranging from 1 to 5, where 1 is the least intense and 5 is the strongest.
Video Number Sadness Relax Happiness Surprise Fear Angry Disgust Neutral
1
2
3
4
5
6
7
8
9
10
11
12
13
14
15
16
Appendix C Interview Questions
  1. 1.
    What color do you like the most? Why?
  2. 2.
    Where is your hometown? Please briefly introduce it.
  3. 3.
    Do you have any hobbies?
  4. 4.
    Have you traveled in the past year?
  5. 5.
    How do you like Beijing?
  6. 6.
    What is your happiest experience?
  7. 7.
    What is your favorite food?
  8. 8.
    What is your personality like?
  9. 9.
    What is your biggest weakness?
  10. 10.
    What is your greatest strength?
  11. 11.
    What do you usually do to relax?
  12. 12.
    Which exercise or sport do you like?
  13. 13.
    Briefly introduce your family members.
  14. 14.
    Who is the person you have the greatest influence on you?
  15. 15.
    Do you have any special places or tourist destinations you want to go to?
  16. 16.
    Who is your favorite celebrity or great person?
  17. 17.
    What is your opinion on the words "neijuan" and "tangping"?
  18. 18.
    What is your favorite literary and artistic work?
  19. 19.
    Have you ever received any rewards or honors in school or at work?
  20. 20.
    What was your most unforgettable experience in the past year?
  21. 21.
    Have you participated in any major event?
  22. 22.
    Have you ever cheated in school or work?
  23. 23.
    Have you concealed a fact to your family or friends in the past year?
  24. 24.
    Have you ever lied to avoid responsibility?